OpenStudy (anonymous):

A regular octagon is inscribed in a circle of radius r. Show that the area of the octagon is 2 root 2r^2

OpenStudy (anonymous):

Use the law of sines to find the area of each of the 8 triangles that form the octagon. Multiply by 8 and you are done.

OpenStudy (anonymous):

\[ \frac{8}{2} r^2 \sin \left(\frac{2 \pi }{8}\right)=4 r^2 \sin \left(\frac{\pi }{4}\right)=\frac{4 \sqrt{2} r^2}{2}=2 \sqrt{2} r^2 \]